OpenStudy (anonymous):

factor by grouping. x^3-2x^2-9x+18

OpenStudy (anonymous):

= x^2(x - 2) - 9(x - 2) = (x^2 - 9)(x - 2) = (x - 2)(x + 2)(x - 2)